if S=-16t^2 + 80t represents the height of a rocket, in feet, t seconds after it was fired, when will the rocket hit the ground? Hint: the rocket is in the ground when s=0
Answer by nerdybill(7384)

if S=-16t^2 + 80t represents the height of a rocket, in feet, t seconds after it was fired, when will the rocket hit the ground? Hint: the rocket is in the ground when s=0
.
start with:
S=-16t^2 + 80t
set s to zero and solve for t:
0=-16t^2 + 80t
0 = -t^2 + 5t
0 = t^2 - 5t
0 = t(t-5)
.
t = {0, 5}
answer: 5 seconds